An investment offers a total return of 15 percent over the coming year. Alex Hamilton thinks the total real return on this investment will be only 77 percent. What does Alex believe the inflation rate will be over the next year? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

Inflation rate = _______ %
Answer #1

EXACT INFLATION RATE=(1+total return)/(1+real return)-1=1.15/1.077-1=6.78%

APPROXIMATE INFLATION RATE=total return-real return=15%-7.7%=7.30%